OpenStudy (anonymous):

convert r=4 cos theta - sin theta to rectangular form

OpenStudy (anonymous):

in equation form....\[r=4\cos \theta - \sin \theta\]

OpenStudy (anonymous):

between Polar and Rectangular coordinate system, we have the ff. relationships: (1) r = sqrt(x^2+y^2) (2) cos theta = x/r (3) sin theta = y/r substitute above relationships to above problem will lead to do the conversion... :-)